OpenStudy (kpop4life123):

What is the length of the missing leg? Round your answer to the nearest tenth.

OpenStudy (kpop4life123):



OpenStudy (anonymous):

Is this a right triangle? If so then you can use Pythagoras theorem, which says: \[a^2+b^2=c^2\]

OpenStudy (anonymous):

\[\sqrt{a^2+b^2}=c\]

OpenStudy (anonymous):

In this case you have a=6 and b=30